OpenStudy (bbb911):

(3/4 - 1)(1/8 + 1/2)

OpenStudy (anonymous):

=(3/4 - 4/4)(1/8 + 4/8) =(-1/4)(5/8) =-5/32

OpenStudy (bbb911):

thank you so much :)